A delivery company allows packages to weigh a maximum of 52 pounds. You packed a box that can still weigh 17 more pounds. Write an equation to find how many pounds your box weighs.
Answer by fractalier(6550)

Call its weight, x. Then we write
x + 17 = 52 (It could be written as an inequality too --> x + 17 <= 52)
Now we subtract 17 from both sides and get
x = 35 pounds